Grade 11Physical Chemistry

A ball of mass 100 g is moving with 100 ms-1. Find its wavelength. Find out atomic number and mass number of the element X. Also identify the element.

Simran Bhatia
12 Years agoGrade 11

1 Answer

Aditi Chauhan
12 Years ago
λ = h/mu = 6.627 *10-34/0.1 *100
or λ = 6.627 * 10-35 m = 6.627 * 10-25